# On the Lebesgue Constant of Extended-Domain Spectral Methods for Elliptic PDEs

### v2

@@ Acknowledgement
For completeness, we provide a proof of the refined lemma concerning the exponential decay of the inverse of a matrix with exponentially decaying entries. The proof is an adaptation of the argument presented by Jaffard for infinite-dimensional operators on $\ell^{2}(\mathbb{Z})$ to the case of a sequence of finite-dimensional matrices [17, Prop. 2].
@@
Let $\{\mathbf{M}_{N}\}_{N\in\mathbb{N}}$ be a sequence of invertible $N\times N$ matrices. Suppose there exist constants $C_{M}>0$ and $\beta>0$, independent of $N$, such that the entries of each $\mathbf{M}_{N}$ satisfy
[…]
@@
$\sup_{N\in\mathbb{N}}\left\lVert\mathbf{M}_{N}^{-1}\right\rVert_{\infty}\leq C_{inv}.$
@@
The proof proceeds in five steps. We first establish the necessary operator-norm bounds, then reduce the problem to the positive definite case, and finally apply the core argument from Jaffard [17].
  1. 1.
@@
    Framework and Uniform Boundedness on $\ell^{2}$: Jaffard defines the matrix algebra $\mathcal{E}_{\gamma}$ for matrices whose entries decay as $e^{-\gamma^{\prime}d(s,t)}$ for all $\gamma^{\prime}<\gamma$ [17, Def. 1]. The hypothesis on $\{\mathbf{M}_{N}\}$ means the sequence is uniformly contained in $\mathcal{E}_{\beta}$.
@@
    By Schur’s Lemma, an operator is bounded on $\ell^{2}$ if its $\ell_{1}$ and $\ell_{\infty}$ induced norms are bounded. The uniform exponential decay of the entries of $\mathbf{M}_{N}$ implies that the sequence is uniformly bounded in both norms:
    $\left\lVert\mathbf{M}_{N}\right\rVert_{\infty}=\max_{i}\sum_{j=1}^{N}|(\mathbf{M}_{N})_{ij}|\leq C_{M}\sum_{k=-\infty}^{\infty}e^{-\beta|k|}=C_{M}\left(1+\frac{2e^{-\beta}}{1-e^{-\beta}}\right)<\infty.$
@@
    The same bound holds for $\left\lVert\mathbf{M}_{N}\right\rVert_{1}$. Consequently, the $\ell_{2}$-norm is also uniformly bounded: $\left\lVert\mathbf{M}_{N}\right\rVert_{2}\leq\sqrt{\left\lVert\mathbf{M}_{N}\right\rVert_{1}\left\lVert\mathbf{M}_{N}\right\rVert_{\infty}}\leq\text{const}$.
@@
    The proof in [17] assumes invertibility on $\ell^{2}$. The hypothesis $\sup_{N}\left\lVert\mathbf{M}_{N}^{-1}\right\rVert_{\infty}\leq C_{inv}$ implies that the sequence of operators is uniformly invertible on $\ell^{2}$. This is because any operator bounded on $\ell_{\infty}$ is also bounded on $\ell^{2}$, establishing the necessary condition to apply the logic of the proof.
  2. 2.
@@
    Reduction to the Positive Definite Case: Following Jaffard [17, p. 464], we analyze the inverse by considering the symmetric positive definite matrix $\mathbf{A}_{N}=\mathbf{M}_{N}\mathbf{M}_{N}^{*}$. The inverse of $\mathbf{M}_{N}$ is then recovered via the identity $\mathbf{M}_{N}^{-1}=\mathbf{M}_{N}^{*}\mathbf{A}_{N}^{-1}$. Since $\mathcal{E}_{\beta}$ is an algebra [17, Prop. 1], and $\{\mathbf{M}_{N}\},\{\mathbf{M}_{N}^{*}\}$ are uniformly in $\mathcal{E}_{\beta}$, their product $\{\mathbf{A}_{N}\}$ is uniformly in $\mathcal{E}_{\beta^{\prime}}$ for any $\beta^{\prime}<\beta$. Furthermore, since $\{\mathbf{M}_{N}\}$ is uniformly invertible on $\ell^{2}$, so is $\{\mathbf{A}_{N}\}$. Our goal is to prove that the entries of $\mathbf{A}_{N}^{-1}$ exhibit uniform exponential decay.
  3. 3.
@@
    Neumann Series Argument: Let $\mathbf{A}$ be a matrix in the sequence $\{\mathbf{A}_{N}\}$. We can write it as $\mathbf{A}=\left\lVert\mathbf{A}\right\rVert_{2}(\mathbf{I}-\mathbf{R})$, where $\mathbf{R}=\mathbf{I}-\mathbf{A}/\left\lVert\mathbf{A}\right\rVert_{2}$. The uniform invertibility of $\{\mathbf{A}_{N}\}$ implies that there exists a constant $r<1$, independent of $N$, such that $\left\lVert\mathbf{R}\right\rVert_{2}\leq r$. The inverse is given by the Neumann series [17, p. 464]:
    $\mathbf{A}^{-1}=\frac{1}{\left\lVert\mathbf{A}\right\rVert_{2}}\sum_{k=0}^{\infty}\mathbf{R}^{k}.$
    The entries of $\mathbf{R}^{k}$ can be bounded in two ways:
@@
      A norm-based bound: $|(\mathbf{R}^{k})_{ij}|\leq\left\lVert\mathbf{R}^{k}\right\rVert_{2}\leq\left\lVert\mathbf{R}\right\rVert_{2}^{k}\leq r^{k}$.
@@
      An entry-wise decay bound: Since $\mathbf{R}$ is in $\mathcal{E}_{\beta^{\prime}}$, repeated application of the algebra property shows that for any $\alpha<\beta^{\prime}$, there is a constant $K>1$ (uniform in $N$) such that $|(\mathbf{R}^{k})_{ij}|\leq K^{k}e^{-\alpha|i-j|}$.
  4. 4.
    Combining Bounds and Summing the Series: Following Jaffard [17, p. 470], we combine these bounds. For any integer $k_{0}>0$, we split the sum:
    $|(\mathbf{A}^{-1})_{ij}|\leq\frac{1}{\left\lVert\mathbf{A}\right\rVert_{2}}\left(\sum_{k=0}^{k_{0}}|(\mathbf{R}^{k})_{ij}|+\sum_{k=k_{0}+1}^{\infty}|(\mathbf{R}^{k})_{ij}|\right).$
@@
    $|(\mathbf{A}^{-1})_{ij}|\leq\text{const}\cdot\left(\sum_{k=0}^{k_{0}}K^{k}e^{-\alpha|i-j|}+\sum_{k=k_{0}+1}^{\infty}r^{k}\right)$
@@
    where $\alpha^{\prime}=\frac{\alpha|\ln(r)|}{\ln(K)-\ln(r)}>0$. The first term exhibits similar behavior. Thus, there exist constants $C^{\prime}>0$ and $\alpha^{\prime}>0$, independent of $N$, such that $|(\mathbf{A}_{N}^{-1})_{ij}|\leq C^{\prime}e^{-\alpha^{\prime}|i-j|}$.
  5. 5.
    Conclusion: We have shown that $\{\mathbf{A}_{N}^{-1}\}$ is uniformly in $\mathcal{E}_{\alpha^{\prime}}$. Since $\{\mathbf{M}_{N}^{*}\}$ is uniformly in $\mathcal{E}_{\beta}$, and $\mathcal{E}$ is an algebra, their product $\mathbf{M}_{N}^{-1}=\mathbf{M}_{N}^{*}\mathbf{A}_{N}^{-1}$ must be uniformly in $\mathcal{E}_{\min(\beta,\alpha^{\prime})}$. This establishes the existence of uniform constants $C$ and $\alpha$ for the decay of entries in $\mathbf{M}_{N}^{-1}$, completing the proof.
@@ Appendix B Technical Lemmas for the Proof of the Maximum Principle
This is the Gårding inequality with the identified constants $\alpha_{0}=1$ and $C_{0}=\frac{1}{2}\left\lVert\nabla\cdot\mathbf{k}\right\rVert_{L^{\infty}}$. ∎
@@
    $H^{1}$-orthogonality defect for the Laplacian: $|\langle\nabla f,\nabla(f-\mathcal{I}_{N}f)\rangle|\leq C_{O}N^{-1}\left\lVert f\right\rVert_{H^{1}}^{2}$. This is a known, non-trivial result from spectral approximation theory (see e.g. [19, Sec. 4].
@@
3. Final Assembly. Summing the bounds for all error components ($E_{1}$ and the parts of $E_{2}$), we find that for any chosen $\epsilon>0$, we can choose $N$ large enough such that the total contribution to the $\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}$ term is less than $\epsilon$. All other terms can be collected into a single term $C_{\epsilon}\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}$, where $C_{\epsilon}$ depends on $\epsilon$, operator norms, and interpolation constants, but is independent of $v_{N}$ and $N$. This completes the proof. ∎

### v3

@@ Acknowledgement
For completeness, we provide a proof of the refined lemma concerning the exponential decay of the inverse of a matrix with exponentially decaying entries. The proof is an adaptation of the argument presented by Jaffard for infinite-dimensional operators on $\ell^{2}(\mathbb{Z})$ to the case of a sequence of finite-dimensional matrices [jaffard1990proprietes, Prop. 2].
@@
Let $\{\bm{M}_{N}\}_{N\in\mathbb{N}}$ be a sequence of invertible $N\times N$ matrices. Suppose there exist constants $C_{M}>0$ and $\beta>0$, independent of $N$, such that the entries of each $\bm{M}_{N}$ satisfy
$|(\bm{M}_{N})_{ij}|\leq C_{M}e^{-\beta|i-j|}\quad\text{for all }1\leq i,j\leq N.$
@@
$\sup_{N\in\mathbb{N}}\left\lVert\bm{M}_{N}^{-1}\right\rVert_{\infty}\leq C_{inv}.$
@@
The proof proceeds in five steps. We first establish the necessary operator-norm bounds, then reduce the problem to the positive definite case, and finally apply the core argument from Jaffard [jaffard1990proprietes].
  1. (1)
@@
    Framework and Uniform Boundedness on $\ell^{2}$: Jaffard defines the matrix algebra $\mathcal{E}_{\gamma}$ for matrices whose entries decay as $e^{-\gamma^{\prime}d(s,t)}$ for all $\gamma^{\prime}<\gamma$ [jaffard1990proprietes, Def. 1]. The hypothesis on $\{\bm{M}_{N}\}$ means the sequence is uniformly contained in $\mathcal{E}_{\beta}$.
@@
    By Schur’s Lemma, an operator is bounded on $\ell^{2}$ if its $\ell_{1}$ and $\ell_{\infty}$ induced norms are bounded. The uniform exponential decay of the entries of $\bm{M}_{N}$ implies that the sequence is uniformly bounded in both norms:
    $\left\lVert\bm{M}_{N}\right\rVert_{\infty}=\max_{i}\sum_{j=1}^{N}|(\bm{M}_{N})_{ij}|\leq C_{M}\sum_{k=-\infty}^{\infty}e^{-\beta|k|}=C_{M}\left(1+\frac{2e^{-\beta}}{1-e^{-\beta}}\right)<\infty.$
@@
    The same bound holds for $\left\lVert\bm{M}_{N}\right\rVert_{1}$. Consequently, the $\ell_{2}$-norm is also uniformly bounded: $\left\lVert\bm{M}_{N}\right\rVert_{2}\leq\sqrt{\left\lVert\bm{M}_{N}\right\rVert_{1}\left\lVert\bm{M}_{N}\right\rVert_{\infty}}\leq\text{const}$.
@@
    The proof in [jaffard1990proprietes] assumes invertibility on $\ell^{2}$. The hypothesis $\sup_{N}\left\lVert\bm{M}_{N}^{-1}\right\rVert_{\infty}\leq C_{inv}$ implies that the sequence of operators is uniformly invertible on $\ell^{2}$. This is because any operator bounded on $\ell_{\infty}$ is also bounded on $\ell^{2}$, establishing the necessary condition to apply the logic of the proof.
  2. (2)
@@
    Reduction to the Positive Definite Case: Following Jaffard [jaffard1990proprietes, p. 464], we analyze the inverse by considering the symmetric positive definite matrix $\bm{A}_{N}=\bm{M}_{N}\bm{M}_{N}^{*}$. The inverse of $\bm{M}_{N}$ is then recovered via the identity $\bm{M}_{N}^{-1}=\bm{M}_{N}^{*}\bm{A}_{N}^{-1}$. Since $\mathcal{E}_{\beta}$ is an algebra [jaffard1990proprietes, Prop. 1], and $\{\bm{M}_{N}\},\{\bm{M}_{N}^{*}\}$ are uniformly in $\mathcal{E}_{\beta}$, their product $\{\bm{A}_{N}\}$ is uniformly in $\mathcal{E}_{\beta^{\prime}}$ for any $\beta^{\prime}<\beta$. Furthermore, since $\{\bm{M}_{N}\}$ is uniformly invertible on $\ell^{2}$, so is $\{\bm{A}_{N}\}$. Our goal is to prove that the entries of $\bm{A}_{N}^{-1}$ exhibit uniform exponential decay.
  3. (3)
@@
    Neumann Series Argument: Let $\bm{A}$ be a matrix in the sequence $\{\bm{A}_{N}\}$. We can write it as $\bm{A}=\left\lVert\bm{A}\right\rVert_{2}(\bm{I}-\bm{R})$, where $\bm{R}=\bm{I}-\bm{A}/\left\lVert\bm{A}\right\rVert_{2}$. The uniform invertibility of $\{\bm{A}_{N}\}$ implies that there exists a constant $r<1$, independent of $N$, such that $\left\lVert\bm{R}\right\rVert_{2}\leq r$. The inverse is given by the Neumann series [jaffard1990proprietes, p. 464]:
    $\bm{A}^{-1}=\frac{1}{\left\lVert\bm{A}\right\rVert_{2}}\sum_{k=0}^{\infty}\bm{R}^{k}.$
    The entries of $\bm{R}^{k}$ can be bounded in two ways:
@@
      A norm-based bound: $|(\bm{R}^{k})_{ij}|\leq\left\lVert\bm{R}^{k}\right\rVert_{2}\leq\left\lVert\bm{R}\right\rVert_{2}^{k}\leq r^{k}$.
@@
      An entry-wise decay bound: Since $\bm{R}$ is in $\mathcal{E}_{\beta^{\prime}}$, repeated application of the algebra property shows that for any $\alpha<\beta^{\prime}$, there is a constant $K>1$ (uniform in $N$) such that $|(\bm{R}^{k})_{ij}|\leq K^{k}e^{-\alpha|i-j|}$.
  4. (4)
    Combining Bounds and Summing the Series: Following Jaffard [jaffard1990proprietes, p. 470], we combine these bounds. For any integer $k_{0}>0$, we split the sum:
    $|(\bm{A}^{-1})_{ij}|\leq\frac{1}{\left\lVert\bm{A}\right\rVert_{2}}\left(\sum_{k=0}^{k_{0}}|(\bm{R}^{k})_{ij}|+\sum_{k=k_{0}+1}^{\infty}|(\bm{R}^{k})_{ij}|\right).$
@@
    $|(\bm{A}^{-1})_{ij}|\leq\text{const}\cdot\left(\sum_{k=0}^{k_{0}}K^{k}e^{-\alpha|i-j|}+\sum_{k=k_{0}+1}^{\infty}r^{k}\right)$
@@
    where $\alpha^{\prime}=\frac{\alpha|\ln(r)|}{\ln(K)-\ln(r)}>0$. The first term exhibits similar behavior. Thus, there exist constants $C^{\prime}>0$ and $\alpha^{\prime}>0$, independent of $N$, such that $|(\bm{A}_{N}^{-1})_{ij}|\leq C^{\prime}e^{-\alpha^{\prime}|i-j|}$.
  5. (5)
    Conclusion: We have shown that $\{\bm{A}_{N}^{-1}\}$ is uniformly in $\mathcal{E}_{\alpha^{\prime}}$. Since $\{\bm{M}_{N}^{*}\}$ is uniformly in $\mathcal{E}_{\beta}$, and $\mathcal{E}$ is an algebra, their product $\bm{M}_{N}^{-1}=\bm{M}_{N}^{*}\bm{A}_{N}^{-1}$ must be uniformly in $\mathcal{E}_{\min(\beta,\alpha^{\prime})}$. This establishes the existence of uniform constants $C$ and $\alpha$ for the decay of entries in $\bm{M}_{N}^{-1}$, completing the proof.
Q.E.D.
@@ Appendix B Technical Lemmas for the Proof of the Maximum Principle
This is the Gårding inequality with the identified constants $\alpha_{0}=1$ and $C_{0}=\frac{1}{2}\left\lVert\nabla\cdot\mathbf{k}\right\rVert_{L^{\infty}}$. Q.E.D.
@@
    $H^{1}$-orthogonality defect for the Laplacian: $|\langle\nabla f,\nabla(f-\mathcal{I}_{N}f)\rangle|\leq C_{O}N^{-1}\left\lVert f\right\rVert_{H^{1}}^{2}$. This is a known, non-trivial result from spectral approximation theory (see e.g. [bernardi1989properties, Sec. 4]).
@@
3. Final Assembly. Summing the bounds for all error components ($E_{1}$ and the parts of $E_{2}$), we find that for any chosen $\epsilon>0$, we can choose $N$ large enough such that the total contribution to the $\left\lVert\nabla v_{N}^{-}\right\rVert_{L^{2}}^{2}$ term is less than $\epsilon$. All other terms can be collected into a single term $C_{\epsilon}\left\lVert v_{N}^{-}\right\rVert_{L^{2}}^{2}$, where $C_{\epsilon}$ depends on $\epsilon$, operator norms, and interpolation constants, but is independent of $v_{N}$ and $N$. This completes the proof. Q.E.D.